Let x be a random variable that represents the percentage of successful free throws a professional basketball player makes in a season. Let y be a random variable that represents the percentage of successful field goals a professional basketball player makes in a season.


The table below shows the percentage of successful free throws and field goals a professional basketball player makes in a season:


x

52

70

76

77

77

86


y

74

56

51

74

72

68


Suppose you want to predict the percentage of successful field goals a professional basketball player using the percentage of successful free throws, what is the value of a? Round your answers to the nearest hundredths.


1
Expert's answer
2022-05-25T14:31:43-0400

In order to compute the regression coefficients, the following table needs to be used:



"\def\arraystretch{1.5}
 \begin{array}{c:c:c:c:c:c}
 & X & Y & XY & X^2 & Y^2 \\ \hline
 & 52 & 74 & 3848 & 2704 & 5476 \\
 \hdashline
 & 70 & 56 & 3920 & 4900 & 3136 \\
 \hdashline
 & 76 & 51 & 3876 & 5776 & 2601 \\
 \hdashline
 & 77 & 74 & 5698 & 5929 & 5476 \\
 \hdashline
 & 77 & 72 & 5544 & 5929 & 5184 \\
 \hdashline
 & 86 & 68 & 5848 & 7396 & 4624 \\
 \hdashline
Sum= & 438 & 395 & 28734 & 32634 & 26497 \\
 \hdashline
\end{array}""\bar{X}=\dfrac{1}{n}\sum _{i}X_i=\dfrac{438}{6}""=73""\bar{Y}=\dfrac{1}{n}\sum _{i}Y_i=\dfrac{395}{6}""=65.833333""SS_{XX}=\sum_iX_i^2-\dfrac{1}{n}(\sum _{i}X_i)^2""=32634-\dfrac{438^2}{6}=660""SS_{YY}=\sum_iY_i^2-\dfrac{1}{n}(\sum _{i}Y_i)^2""=26497-\dfrac{395^2}{6}=492.833333"




"SS_{XY}=\sum_iX_iY_i-\dfrac{1}{n}(\sum _{i}X_i)(\sum _{i}Y_i)""=28734-\dfrac{438(395)}{6}=−101""b=\dfrac{SS_{XY}}{SS_{XX}}=\dfrac{-101}{660}""=-0.15""a=\bar{Y}-b\bar{X}""a=\dfrac{395}{6}-(\dfrac{-101}{660})(73)=77.00"


"Y=77.00\u22120.15X"